OpenStudy (anonymous):

The base length of a particular triangle is 8 meters more than twice its height. A new triangle is formed by increasing the base by 10 meters and increasing the height by 5 meters. The area is 65 square meters more than the area of the original triangle. Find the base length and height of the original triangle.

OpenStudy (anonymous):

Can you show me your approach ??

OpenStudy (anonymous):

I am confused on this problem.. I have tried to do it but I keep getting confused

OpenStudy (anonymous):

let the base of triangle1 be "b" and height be "h" => b = 8+2h A1 = b*h/2 = h*(h+4) (after simplification) Can you try the same thing with triangle2 ??

OpenStudy (anonymous):

I was thinking triangle 2 would be like , b + 10 and h + 5. That last part of what you wrote is confusing to me.

OpenStudy (anonymous):

i just found the area of the first triangle... A1=(8+2h)*h/2 = 2*(h+4)*h/2=(h+4)*h

OpenStudy (anonymous):

for triangle2 A2 = (b+10)(h+5)/2 = (18+2h)*(h+5)/2 = (h+5)*(h+9) Also A2 = A1 + 65

OpenStudy (anonymous):

Hope that helps